OpenStudy (dancergirl45):

NEED HELP FAST WILL FAN AND MEDAL Solve for x: 5|2x − 2| + 8 = 18

OpenStudy (dancergirl45):

@TrojanPoem

OpenStudy (igigighjkl):

@mathmale

OpenStudy (trojanpoem):

|2x-2| can be postive or negative so 5(2x-2) + 8 = 18 or 5(-2x+2) + 8 = 18 solve for x in both cases.

OpenStudy (trojanpoem):

10x - 10 + 8 = 18 10x - 2 = 18 10x = 20 x = 2 -10x + 10 + 8 = 18 -10x = 0 x = 0 Solution set = {0,2}

OpenStudy (dancergirl45):

Ok thank you so much!!